OpenStudy (anonymous):

Circle A and circle B are congruent. CD is a chord of both circles. If CD = 16 cm and the radius is 17 cm, how long is AB

OpenStudy (primeralph):

@Hero

OpenStudy (anonymous):

i got 23 is it right

OpenStudy (primeralph):

@campbell_st Bed time for me.

OpenStudy (anonymous):

17/2=8.5+17=AB

OpenStudy (anonymous):

8.5=8.5+17

OpenStudy (anonymous):

@huneya

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is the answer 23? @huneya

OpenStudy (anonymous):

no its 25.5

OpenStudy (anonymous):

so i will go with 23 since its close to 25.5 which i think is the answer

OpenStudy (anonymous):

kay

hero (hero):

@julietorres1, you could have at least drawn it out. I don't know what it looks like.

OpenStudy (mckenzieandjesus):

@Hero this is what the pic looks like for this question is it 23?

hero (hero):

yep.

hero (hero):

Do you know what to do to solve it?

OpenStudy (mckenzieandjesus):

no :/